OpenStudy (anonymous):

Find S15 for 1 + 1.5 + 2.25 + 3.375 +...? use the general equation Sn= T1((1-r^n)/(1-r)

OpenStudy (anonymous):

I got: r= 1.5/1 = 1.5 ?? cause that equation I was given to find r is r= t1/t2

OpenStudy (anonymous):

and then I put it in as S15=1((1-1.5^15)/(1-5))= 873.79..

OpenStudy (anonymous):

but I don't think this is right.